OpenStudy (anonymous):

betty gets paid 25 cents per paper to deliver newspapers in her neighborhood. when betty is sick, she pays her brother bobby 10 cents per paper to deliever them for her. if bobby makes $7.80 cents filling in for betty, how much money does betty get to keep for that day?

OpenStudy (anonymous):

Bobby makes 7.80 , getting 10 cents per paper Number of papers is 7.80*100/10 = 78 paper For each paper Betty gets 25 -10 = 15 cents For 78 papers Betty will get 78 * 0.15 =$11.7 for that day

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

can you explain it one step at a time please?

OpenStudy (anonymous):

Bobby makes 7.80 He gets 10 cents per paper Total amount he got is $7.80 that is 780 cents Number of papers = 780/ 10 = 78

OpenStudy (anonymous):

For one paper Betty gets 25 cnets, out of this she gives bobby 10 cents. For one paper Betty has 25 -10 = 15 cents

OpenStudy (anonymous):

780 papers, 15 cents for each paper Total = 780 * 0.15 = $11.70

OpenStudy (anonymous):

ohhhhh kk thanks soo much:D

OpenStudy (anonymous):

Welcome